3.2.17 e statlistics class taught unsing traditional lechure and a class taught using a ipped classroom Complete parts (a) theough (c) below The following dala represent exam scores in a Traditional 702 699 795 67 5 865 774 565 825 798 705 64.7 692 59 770 708 633 734 776 924 794 764 833 709 916 770 764 Flipped (e) Which course has more dispersion in exam scores using the range as the measure of dispersion? The tadinal course has a range of B while the neer course has a range ofㄿ Thebdi nal corse has more diyersin Type integers or decimals. Do not round) (bj Which course has more dopersion in exam scores using the sample standard deviation as the measure of dispersion? The tadtinal course has a standard devation of 8993), while tho pped couse has a standard dervilion of 8019Thetraditional course has more dspersion (Round to three decimal places as needed )
Answer #1

Solution:-
(a) Range :

Traditional :- Range is 30

Explanation

The range is the difference between the highest and lowest values in the data set.

Ordering the data from least to greatest, we get:

56.5 59.0 64.7 67.5 69.2 69.9 70.2 70.5 77.4 79.5 79.8 82.5 86.5   

The lowest value is 56.5.

The highest value is 86.5.

The range = 86.5 - 56.5 = 30.

  

Flipped : Range = 29.1

Explanation

The range is the difference between the highest and lowest values in the data set.

Ordering the data from least to greatest, we get:

63.3 70.8 70.9 73.4 76.4 76.4 77.0 77.0 77.6 79.4 83.3 91.6 92.4   

The lowest value is 63.3.

The highest value is 92.4.

The range = 92.4 - 63.3 = 29.1.

The traditional course has a range of 30.while the flipped course has a range of 29.1.the traditional course has more dispersion.

b)

(b) Standard deviation:-

traditional course: 8.975

Explanation To find standard deviation we use the following formula , Σ(n x) We will compute this formula in 4 steps. Step 1: Find the mean. In this example: X 71.7846 (use this calculator for step by step explanation on how to find mean). Step 2: Create the following table. data 70.2 69.9 79.5 87.5 88.5 77.4 56.5 82.5 79.8 70.5 4.7 89.2 59.0 data-mean 1.5846 1.8846 7.7154 4.2848 14.7154 5.8154 -15.2848 10.7154 8.0154 1.2848 7.0848 2.5848 -12.7848 (data - mean) 2.51095716 3.55171716 59.52739718 18.35779718 216.54299716 31.53271718 233.81899718 114.81979718 84.24863718 1.85019718 50.19155716 8.88015716 183.44599718 Step 3: Find the sum of numbers in the last column to get. , x)966.6769 Step 4: Calculate ơ using the above formula. Σ(n x)-/066.6760~8.9753 13-1


Flipped course : 7.998

the traditional course has a standard deviation 8.975, while the flipped course has a standard deviation of 7.998,the traditional course has more dispersion.